### v1

# GVGAI-LLM: Evaluating Large Language Model Agents with Infinite Games

Abstract
We introduce GVGAI-LLM, a video game benchmark for evaluating the reasoning and problem-solving capabilities of large language models (LLMs). Built on the General Video Game AI framework, it features a diverse collection of arcade-style games designed to test a model’s ability to handle tasks that differ from most existing LLM benchmarks. The benchmark leverages a game description language that enables rapid creation of new games and levels, helping to prevent overfitting over time. Each game scene is represented by a compact set of ASCII characters, allowing for efficient processing by language models. GVGAI-LLM defines interpretable metrics, including the meaningful step ratio, step efficiency, and overall score, to assess model behavior. Through zero-shot evaluations across a broad set of games and levels with diverse challenges and skill depth, we reveal persistent limitations of LLMs in spatial reasoning and basic planning. Current models consistently exhibit spatial and logical errors, motivating structured prompting and spatial grounding techniques. While these interventions lead to partial improvements, the benchmark remains very far from solved. GVGAI-LLM provides a reproducible testbed for advancing research on language model capabilities, with a particular emphasis on agentic behavior and contextual reasoning.
Introduction
Large language models (LLMs) have shown impressive capabilities in open-ended generation, planning, and reasoning tasks (Achiam et al. 2023; Liu et al. 2024a). Recently, researchers have begun deploying LLMs as interactive agents in sequential decision making environments such as robotics, web navigation, and games (Achiam et al. 2023; Liu et al. 2024a; Yang, Kleinman, and Harteveld 2025). While benchmarks such as GLUE (Wang et al. 2019b), SuperGLUE (Wang et al. 2019a), and MMLU (Hendrycks et al. 2021) evaluate static language understanding and broad domain knowledge, HumanEval (Chen et al. 2021) assesses code generation proficiency, HELM (Liang et al. 2023) examines safety and robustness, and instruction-following frameworks like ALFWorld (Shridhar et al. 2021) and BabyAI (Chevalier-Boisvert et al. 2019) focus on executing textual commands. None of these benchmarks, however, measure decision making competence in structured symbolic domains with game-style logic, real-time decision, and spatial reasoning.
To address this gap, we introduce GVGAI-LLM, a benchmark that adapts the General Video Game AI (GVGAI) framework (Perez-Liebana et al. 2019; Liébana et al. 2019) into a testbed tailored for LLM agents. Unlike existing benchmarks focused on task completion or instruction following, GVGAI-LLM emphasizes goal-directed agent behavior in reactive, rule-based 2D games. The GVGAI framework’s diverse game dynamics and formal Video Game Description Language (VGDL) specifications (Ebner et al. 2013; Schaul 2013) make it ideal for evaluating how well LLMs can parse structured environments, comply with implicit goals, and act.
GVGAI-LLM is designed to evaluate language-only agents in symbolic environments, without access to code execution, simulators, or structured planners. It provides a unified testbed that supports natural language interfaces, tracks interpretable behaviors, and captures unique LLM-specific challenges such as spatial misalignment and inconsistent rule-following. Our key contributions are as follows:
  • We introduce a standardized prompting interface that enables large language model-based agents to interact with more than one hundred games via natural language, supporting zero-shot and contextual prompting.
  • We reformulate symbolic game states into structured textual representations, making them accessible to language-only agents without reliance on internal simulators or programmatic logic.
  • We design interpretable and reproducible metrics, including meaningful step ratio, step efficiency, and win rate, to assess agent behavior and reveal key failure modes.
Figure 1: Overview of GVGAI-LLM
Related Work
Games as LLM Benchmarks.
Games have recently emerged as powerful benchmarks for evaluating a broad range of LLM capabilities, including reasoning, planning, spatial understanding, and multimodal integration. BALROG (Paglieri et al. 2025) assesses LLM agentic behavior across six environments, such as BabyAI (Chevalier-Boisvert et al. 2019) and Nethack (Küttler et al. 2020). Several other benchmarks, SmartPlay (Wu et al. 2024), AgentBench (Liu et al. 2024b), DSGBench (Tang et al. 2025), GTBench (Duan et al. 2024), and LMGame-Bench (Hu et al. 2025), evaluate decision-making and reasoning through a variety of games. GameTraversalBenchmark (Nasir, James, and Togelius 2024) focuses on planning via navigation in procedurally generated levels, while ARC-AGI (Chollet et al. 2024, 2025) emphasizes abstract reasoning through structured puzzles.
Other benchmarks target multimodal understanding, such as Atari-GPT (Waytowich et al. 2024), VideoGameBench (Zhang et al. 2025), and MMBench (Liu et al. 2024c), which require interpreting visual inputs alongside text. These benchmarks highlight the breadth of skills being evaluated in games, yet many rely on either static environments, pre-defined trajectories, or focus primarily on instruction following or multimodal grounding.
In contrast, GVGAI-LLM emphasizes zero-shot, agent-driven decision-making in reactive, rule-based games through natural language interfaces. By focusing on symbolic logic, spatial dynamics, and prompt-based interaction, it fills a critical gap in evaluating LLM reasoning in structured and discrete environments.
Despite the variety of benchmarks, LLMs generally perform poorly across most game-based tasks, particularly those involving spatial reasoning, long-horizon planning, or symbolic manipulation. Their performance often lags behind traditional planning and reinforcement learning agents, especially in environments with sparse rewards or complex rules. Furthermore, inference with LLMs remains computationally expensive, as token-based prompting and decoding must be repeated at every step of interaction.
Figure 2: VGDL rule mapping and level layout mapping
GVGAI Framework.
The original GVGAI framework (Perez-Liebana et al. 2019; Liébana et al. 2019) provides a unified testbed for evaluating general game-playing agents across diverse environments. It emphasized planning, reactivity, and generalization. Game rules and levels are represented by VGD (Schaul 2013), allowing for the creation of new games and levels to evaluate agents (Khalifa et al. 2016), whereas existing benchmarks typically provide only a single or a fixed set of environments.
GVGAI has been extensively used to evaluate planning agents and reinforcement learning methods, emphasizing generalization, reactivity, and symbolic reasoning (Hu et al. 2022). While it was not originally designed for language-based agents, its modular, rule-based design enables structured interactions that are well-suited for LLM prompting. Modifying GVGAI to accommodate LLM agents enables direct comparison between LLMs, tree search, and RL agents in a shared symbolic space.
Unlike many instruction-focused benchmarks, GVGAI-LLM poses distinct challenges centered on spatial grounding, symbolic dynamics, and reactive decision-making. We provide experimental comparisons across agent types and summarize performance differences in Table 3.
Benchmark Design
Our benchmark is built on top of the original GVGAI Java engine, which offers a collection of over one hundred 2D video games featuring a variety of tasks, challenges, and skill levels, all defined using VGDL, which also enables rapid creation of new games and levels, making the benchmark infinitely extensible. To enable interaction with language models, we substantially extend the GVGAI framework by exposing both the dynamic game state at each time step and the underlying game rules as structured textual representations. This includes translating symbolic state information and formal game mechanics into natural language descriptions suitable for prompting language models. Specifically, each environment is serialized into a symbolic 2D layout, paired with a natural language goal description and a mapping of game entities to their meanings.
Environment and Agent Design
In this subsection, we describe the components responsible for rule translation, prompt formatting, agent decision making, and interaction protocols. The GVGAI environment executes the game logic and provides access to both the symbolic game state and the underlying VGDL rules at each time step. These are processed by the prompt generation module, which converts the state into a structured 2D layout and translates the game rules into natural language. The resulting prompt includes goal descriptions, entity mappings, available actions, and coordinate annotations. This prompt is then passed to the LLM agent, which selects an action based solely on the current step’s information. The selected action is executed in the environment, completing a closed loop. The overall design (c.f., Figure 1) aims to isolate per-step reasoning and encourage symbolic generalization without reliance on temporal memory.
Translator
VGDL (Ebner et al. 2013; Schaul 2013) is a game description language that encodes game rules and levels/scenes as a compact string of ASCII characters. The Translator module converts VGDL rule specifications into natural language descriptions. It parses object definitions, interaction rules, and termination conditions from the VGDL script. For example, a rule such as avatar > key = > killSprite is rendered as: “If the avatar touches a key, the key disappears and the avatar obtains it.” These textual rule descriptions are provided to downstream modules to support language-based reasoning, as shown in Figure 2.
Player
The Player module receives the current game state (represented in ASCII format) along with a high-level strategy. It selects a concrete low-level action (e.g., “move right”) aimed at fulfilling the strategy objective. All decisions are made through language-based reasoning without access to symbolic forward simulation.
Prompt Configuration
The zero-shot prompt includes several static and dynamic components summarized in Table 1. These elements are presented to the agent at each step to facilitate rule interpretation, spatial reasoning, and action selection. Empirically, we observe that each step consumes approximately 8,000 tokens when using reasoner models (e.g., GPT-o3-mini) and around 5,000 tokens for standard models. This includes both static components (such as game rules) and dynamic components (such as current state).
Prompt Component Zero-Shot Setting
Static Rule Description Included (each step)
Action Mapping Included (each step)
Current Game State Included
Sprite Mapping Included
Avatar Position Included
Past States / Actions Not included
Table 1: Prompt components used in the zero-shot setting.
Figure 3: Comprehensive score of each model across six GVGAI games. The comprehensive score aggregates meaningful action ratio, inverse steps, and (when available) reward and win rate.
LLM Agent Interaction Pipeline
GVGAI-LLM follows a simple yet strict interaction loop between the environment and the language model. At each time step, the environment generates a structured textual description of the current state, including an ASCII map, rule explanations, a list of available actions, and the avatar’s current position. This information is formatted as a prompt and sent to the language model, which responds with an action choice in natural language. The environment then applies this action, updates the state, and prepares the next prompt based on the new state.
Importantly, each step is handled independently. The model does not receive any memory of previous states or actions, nor does it accumulate trajectory information. This zero-shot setup ensures that the agent must make decisions based entirely on the current state. This design of the prompt encourages reasoning over symbolic input rather than reliance on historical context or pattern across time.
Contextual Prompting
Although our benchmark adopts a zero-shot setup by default, we also explored a contextual prompting variant in which each input includes a short history, up to 20,000 tokens, comprising prior states, actions, and chat. Since prompts are constructed as interactive dialogues, this incorporates previous exchanges to provide the model with a limited form of temporal continuity. However, we found this approach introduced several drawbacks, including compounded reasoning errors, and higher token costs, without significantly improving task performance.
First, the addition of historical context did not lead to higher win rates or better decision quality across most games. Second, the extra context greatly increased the prompt length, which leads to higher token usage and cost. Finally, including prior steps shifted the focus away from general reasoning and symbolic interpretation, instead encouraging models to rely on memorized patterns. Since GVGAI-LLM is designed to evaluate how well models can generalize from the current state, we chose to standardize on zero-shot prompting as our main evaluation setting.
Evaluation Metrics and Scoring Framework
To assess the behavior of LLM agents in GVGAI-LLM, we introduce a set of evaluation metrics that capture key aspects of decision-making quality.
Meaningful Step Ratio
The first metric is the meaningful step ratio, which quantifies the proportion of agent actions that produce tangible effects on the environment. This metric captures the agent’s ability to engage in purposeful, goal-directed behavior and helps differentiate between effective and redundant actions.
A step is considered meaningful if and only if it satisfies:
$\texttt{isMeaningful(step)}=(\Delta\text{reward}\neq 0)\lor(\Delta\text{state}\neq 0)$
where $\Delta\text{state}\neq 0$ indicates that the action causes a visible change in the game world, such as the removal, transformation, or appearance of an entity. We explicitly exclude the following from meaningful steps:
  • ACTION_NIL (i.e., doing nothing at current time step);
  • Repeating ineffective actions (e.g., walking into a wall);
  • Moves that cancel each other trivially in the last 4 steps (e.g., moving left immediately after moving right).
A higher meaningful step ratio indicates that the agent is frequently executing impactful actions that contribute to progress toward the game’s objectives. This metric is particularly relevant for evaluating reasoning, planning, and the agent’s ability to interact meaningfully with symbolic game environments. Notably, weak agents, regardless of their underlying architecture, often exhibit low meaningful step ratios, as they tend to perform ineffective behaviors such as moving back and forth or repeating trivial actions.
Step Efficiency
The second metric is step efficiency, which evaluates the agent’s ability to achieve objectives with minimal effort. It is computed as the normalized difference between the average number of steps required to win and the maximum tested steps in a level. Formally, we define:
$\text{Step Efficiency}=1-\frac{\text{Average Win Step}}{\text{Max Step}}.$ (1)
This value ranges from 0 to 1, with higher values indicating more efficient solutions.
Win Rate
In addition to these, we also track win rate, defined as the fraction of episodes in which the agent successfully completes the level, and total game reward, normalized to the interval $[0,1]$. While win rate captures binary task completion, normalized reward offers a more granular perspective by measuring partial progress in cases where the agent does not reach the goal.
Normalized Reward
A normalized reward is computed using min-max scaling across all agent runs for a level:
$R=\frac{r-r_{\min}}{r_{\max}-r_{\min}+\epsilon},$
where $r$ is the total reward of a run, $r_{\min}$ and $r_{\max}$ are the minimum and maximum rewards observed across all agents for the same game level, and $\epsilon$ is a small constant to avoid division by zero.
Overall Score
To summarize agent behavior across these dimensions, we define an overall score by averaging four normalized metrics: meaningful step ratio, total steps (inverted), reward, and win rate. We deliberately assign equal weights to each metric to ensure balanced evaluation across different aspects of agent behavior, and to maintain interpretability across diverse games. Exploratory tests with manual reweighting yielded no consistent improvements in ranking robustness. Each component captures a distinct behavioral property, meaningful step ratio, efficiency, and task success, allowing the final score to reflect multiple dimensions of quality. We deliberately avoid assigning manual weights to prevent overfitting to any single behavior type. This formulation supports a robust, interpretable measure for comparing LLM agents across tasks and prompting strategies. Since reward distributions vary significantly across GVGAI games, we apply per-game normalization to ensure that the normalized reward captures relative performance within each environment.
Experiments
We conduct two experiments to evaluate language model agents in GVGAI-LLM under a standardized zero-shot setting. In this setup, agents choose actions based solely on the current symbolic state, without memory, trajectory history, or access to forward simulation. All models are evaluated via a shared API using identical prompt templates, input formatting, and decoding parameters to ensure comparability. The first experiment evaluates GPT-4o-mini across the entire benchmark, while the second compares multiple LLMs on a selected subset of six games.
Full-Benchmark Evaluation with GPT-4o
In the first experiment, we evaluate GPT-4o-mini across all 118 games in the GVGAI-LLM benchmark. Each game is associated with a single rule file and up to five distinct level layouts. For each level, we run one episode using the same underlying game rules, allowing us to assess the model’s ability to generalize across structurally different scenarios within the same game. Results are summarized in Table 2, with full per-game data in the Appendix. GPT-4o-mini fails to complete the vast majority of game levels in GVGAI-LLM, even when the levels are relatively small and conceptually straightforward for humans. These results underscore the challenge posed by GVGAI-LLM and highlight its value in identifying persistent failures in symbolic reasoning and spatial decision-making.
Figure 4: Win rate comparison of three LLM agents across six GVGAI games. The win rate is computed as the percentage of successful completions out of multiple rollouts.
Metric Value
Total Games Tested 118
Levels with 0% Win Rate 477/540
Average Meaningful Step Ratio 49.71%
Average Step Efficiency 0.3293
Overall Win Rate 10.27%
Average Overall Score 0.2764
Best Performing Game (Win Rate) Cec3 (100.00%)
Table 2: GPT-4o summary statistics.
Agent/Model Aliens Boulderdash Escape Realsokoban Sokoban Zelda
DeepSeek-chat 16.7% 0.0% 20.0% 0.0% 0.0% 23.3%
Deepseek-r1 0.0% 0.0% 54.5% 0.0% 50.0% 28.6%
GPT-4o-mini 8.0% 0.0% 4.0% 0.0% 0.0% 0.0%
GPT-o3-mini 80.0% 0.0% 44.0% 0.0% 52.0% 72.0%
Gemini-2.5-pro 16.0% 0.0% 36.0% 4.0% 28.0% 8.0%
gemini-2.0-exp.-flash 4.0% 0.0% 12.5% 0.0% 0.0% 0.0%
olets 100.0% 56.0% 68.0% 0.0% 24.0% 76.0%
sampleMCTS 100.0% 28.0% 0.0% 0.0% 40.0% 24.0%
sampleRHEA 100.0% 4.0% 24.0% 0.0% 24.0% 36.0%
baseline3DQN 56.0% 0.0% 16.0% 0.0% 0.0% 0.0%
baseline3PPO 64.0% 0.0% 12.0% 0.0% 12.0% 8.0%
Table 3: Win rates comparison across planning agents in GVGAI and LLMs.
Multi-Model Comparison on Selected Games
In the second experiment, we compare multiple LLMs on a curated subset of six representative games: zelda, aliens, boulderdash, realsokoban, escape, and sokoban with five levels per game. These games were chosen to span a variety of mechanics and difficulty degrees, ranging from real-time action (e.g., aliens, zelda) to spatial puzzles (sokoban, realsokoban) and open-ended navigation tasks (escape). The selection was informed by procedural content generation (PCG) design taxonomies (Shaker, Togelius, and Nelson 2016; Yannakakis and Togelius 2018), ensuring diversity (Li et al. 2025) in spatiality, interactivity, and temporal structure.
We evaluate six models: gpt-4o-mini, o3-mini, gemini-2.0-flash-exp, gemini-2.5-pro (without thinking), deepseek-chat, and Deepseek-r1. Each model is run under the same zero-shot prompting setup, with five repeated runs per game level. All prompts and decoding parameters are held nearly constant with a temperature of 0.9. Although we explored contextual prompting, we ultimately discarded it due to high token costs and limited performance gains. This experiment focuses on comparing model behavior across games with varied symbolic and spatial demands. Detailed results are shown in Appendix.
Results
The multi-model evaluation reveals clear differences in model behavior and competence across symbolic environments. gpt-4o-mini demonstrates relatively balanced performance, achieving moderate-to-high scores in puzzle-based games like sokoban and realsokoban, although it still fails to consistently complete levels. o3-mini shows similar trends, with stable but limited competence in levels requiring deterministic planning, such as escape and realsokoban, yet with overall low win rates and little generalization across level variations. In contrast,gemini-2.0-flash-exp achieves moderate scores but rarely succeeds in completing any levels, reflecting minimal spatial grounding. The more capable gemini-2.5-pro occasionally solves simple levels, particularly in Escape, but still underperforms on more complex puzzle scenarios. Meanwhile, deepseek-chat achieves a few wins in early levels but exhibits erratic behavior and low overall score across most puzzle-based games.
The results show that language model agents perform inconsistently across games and generally underperform compared to search-based methods. GPT-o3-mini stands out among LLMs with relatively strong performance on games like Aliens and Zelda, but still falls short of search-based agents in overall reliability. Reinforcement learning baselines perform poorly overall. baseline3DQN succeeds only in Aliens (56.0%) and Escape (16.0%), while baseline3PPO performs slightly better in Aliens (64.0%) and achieves low win rates in Escape (12.0%), Sokoban (12.0%), and Zelda (8.0%). Their near-zero performance in planning-heavy games highlights the challenge posed by sparse rewards and long-horizon reasoning in GVGAI-LLM. However, in planning-heavy environments such as Sokoban and Escape, some LLMs show surprising strengths. For instance, Deepseek-r1 achieves 50.0% in Sokoban and 54.5% in Escape, outperforming many other models. This suggests that LLMs may possess useful priors or structured reasoning capabilities that help in environments where search alone is not sufficient.
Discussion
GVGAI-LLM reveals systematic limitations in how current language-only models reason about symbolic, spatially grounded environments. Across diverse games and agent types, we observe consistent failure patterns that cannot be explained by randomness or decoding noise alone. These behaviors suggest deeper challenges in symbolic understanding of the level, spatial inference, and action reasoning. We group our findings into three core categories: (1) spatial grounding failures, (2) symbolic identity confusion, and (3) behavioral misalignment.
Spatial Grounding Errors
Despite receiving structured textual prompts, LLMs frequently misinterpret spatial layouts. First, coordinate confusion, where the model reverses vertical orientation or misaligns row and column positions, often leads to incorrect assumptions about adjacency. Second, hallucinated proximity in sparse layouts causes agents to misjudge distances (e.g., assuming two far-apart objects are adjacent due to horizontal alignment in ASCII maps). These failures often result in ineffective movement and wandering behavior.
We address this issue by introducing two prompt modifications. Explicit coordinate tagging removes ambiguity by anchoring object positions numerically (e.g., row=3, col=4). Verbose grounding transforms spatial context into a set of declarative facts (e.g., “row 2, col 7 is a”). These changes improve LLM understanding across games but do not fully resolve the underlying difficulties.
More fundamentally, we observe that LLMs lack the capacity for path planning in the algorithmic sense. Unlike traditional methods such as A, which search over possible trajectories to compute optimal paths, LLMs operate with limited forward context and no internal simulation of future states. While a reasoning-aware model might theoretically learn to emulate planning through intermediate steps, our outputs do not reveal such structured traces. This limitation aligns with findings from GameTraversalBenchmark (Nasir, James, and Togelius 2024), where LLM agents similarly failed to perform multi-step navigation tasks.
Symbolic Identity Confusion
A second class of failures concerns the model’s ability to track symbolic identity over time. In many GVGAI games, entity transformations occur through rule-based interactions, for example, an avatar picking up a key transitions from nokey to withkey. Despite prompts including both the rule and the updated symbol mapping, some models, such as Gemini-pro, treat withkey as a new and unrelated entity. They often respond as if the avatar has disappeared or cannot act. This failure is not due to prompt ambiguity, as the same information is handled correctly by other models (e.g., GPT-o3-mini), indicating a model-internal reasoning limitation.
Rather than correcting for this behavior via external prompting, GVGAI-LLM treats symbolic identity tracking as an intentional test dimension. The benchmark aims to assess whether LLM agents can maintain referential continuity across symbolic transitions, which is fundamental to robust reasoning in structured environments.
Behavioral Misalignment
A third failure class involves misalignment between observed state and chosen action. The most common case is the inappropriate use of ACTION_NIL, i.e., choosing to do nothing even when meaningful interactions are possible.
In several benchmark games, agents encounter interactive objects such as keys, switches, or portals, yet frequently choose to remain stationary. This behavior reveals a consistent failure to act, even when progression clearly requires explicit movement or interaction. While the underlying cause is uncertain, we consider several possible explanations. One possibility is that models assume proximity alone is sufficient to trigger effects, misinterpreting the environment as real-time or event-driven (such as Wait for Breakfast). Another is that standing still acts as a default behavior when the model is uncertain, reflecting indecision rather than a specific belief about the environment. These cases suggest a broader misunderstanding of game dynamics and goal-directed behavior, ultimately limiting task completion.
Such behavioral misalignment persists even when prompts explicitly describe win conditions and available actions. This suggests that the model may struggle to translate symbolic state representations into effective actions. A common failure mode involves incorrect spatial reasoning: for example, the agent may continuously walk into a wall that separates it from a key, seemingly aware of the goal but unable to plan a valid path around the obstacle. Similarly, the agent may fail to recognize that a box can be pushed to reach a target location. These patterns indicate that the model often lacks task-specific associations between objects and actions, or cannot reason over multi-step interactions.
Prompt Design Implications
The above failure modes emphasize the importance of structured prompt design in symbolic, spatial decision environments. Our findings show that LLMs benefit from explicitly encoding spatial relationships, but still struggle with generalization and abstraction in complex layouts.
To address spatial reasoning issues, GVGAI-LLM introduces two prompt-level mitigation strategies: (1) explicit coordinate tagging, which anchors entities to precise locations, and (2) verbose spatial grounding, which guides the model to explicitly assess what lies in each adjacent direction, helping it reason about possible movement options based on local surroundings rather than relying solely on global map structure.
Conclusion
We introduce GVGAI-LLM, a benchmark that reuses and extends the classic GVGAI environment for LLM agents. The benchmark uses textual interfaces, defines interpretable metrics, and ensures reproducibility. It offers a compact setting for analyzing how language models make decisions in games. Our experiments show that most models fail on most games, even when the environments are small or conceptually clear. This reveals fundamental weaknesses in symbolic reasoning, spatial understanding, and planning. GVGAI-LLM remains a difficult challenge and can support further study of LLM behavior in structured decision problems. In the future, we plan to extend GVGAI-LLM to support language models that not only play games, but also design them. This includes using LLMs to generate new game rules and levels, which can then serve as novel test cases for evaluating other models, shifting the focus from agentic behavior to creative and generative capabilities. These design tasks may also provide a new lens on how LLMs reason about rules and game mechanisms, rather than just executing them.
Appendix
A. ChatGPT-4o-mini on more games
Game M.Ratio Win% Score Eff.
Aliens 51.8 0.0 0.213 0.705
Angelsdemons 69.8 0.0 0.285 0.857
Assemblyline 56.1 0.0 0.364 0.137
Avoidgeorge 77.9 0.0 0.257 0.755
Bait 9.6 12.5 0.133 0.000
Beltmanager 48.5 0.0 0.174 0.611
Blacksmoke 82.1 0.0 0.277 0.136
Boloadventures 48.7 0.0 0.326 0.633
Bomber 40.2 0.0 0.155 0.114
Bomberman 86.4 0.0 0.255 0.000
Boulderchase 51.7 0.0 0.185 0.175
Boulderdash 19.9 0.0 0.213 0.159
Brainman 62.9 0.0 0.450 0.476
Bravekeeper 72.3 0.0 0.242 0.827
Butterflies 76.4 80.0 0.512 0.791
Cakybaky 72.4 0.0 0.225 0.781
Camelrace 53.9 20.0 0.232 0.000
Catapults 0.0 0.0 0.040 0.000
Cec1 31.6 0.0 0.117 0.174
Cec2 23.3 0.0 0.219 0.188
Cec3 58.7 100.0 0.443 0.860
Chainreaction 67.0 0.0 0.244 0.844
Chase 55.9 0.0 0.391 0.776
Chipschallenge 58.7 0.0 0.236 0.801
Chopper 79.1 0.0 0.264 0.787
Clusters 72.8 0.0 0.227 0.860
Colourescape 3.7 0.0 0.157 0.000
Cookmepasta 31.7 0.0 0.366 0.478
Cops 47.0 0.0 0.391 0.665
Crossfire 82.2 0.0 0.297 0.644
Decepticoins 50.1 60.0 0.320 0.809
Deceptizelda 32.6 100.0 0.374 0.000
Defem 62.5 0.0 0.199 0.980
Defender 79.4 60.0 0.461 0.728
Deflection 0.0 0.0 0.037 0.510
Digdug 32.2 0.0 0.194 0.175
Donkeykong 59.6 0.0 0.266 0.045
Doorkoban 40.7 0.0 0.263 0.630
Dungeon 78.6 0.0 0.390 0.892
Eggomania 57.5 0.0 0.190 0.739
Eighthpassenger 66.9 0.0 0.219 0.000
Enemycitadel 68.9 0.0 0.416 0.525
Escape 44.9 0.0 0.179 0.566
Factorymanager 47.9 100.0 0.504 0.238
Firecaster 14.2 0.0 0.324 0.136
Fireman 55.8 0.0 0.254 0.160
Firestorms 72.4 0.0 0.294 0.737
Flower 71.6 100.0 0.868 0.980
Freeway 90.6 0.0 0.275 0.000
Frogs 52.7 0.0 0.173 0.493
Greedymouse 14.5 0.0 0.080 0.081
Gymkhana 19.4 0.0 0.089 0.000
Hungrybirds 27.9 0.0 0.209 0.463
Iceandfire 38.4 0.0 0.156 0.602
Ikaruga 62.4 0.0 0.206 0.000
Infection 96.3 80.0 0.607 0.000
Intersection 78.1 0.0 0.251 0.867
Invest 60.5 0.0 0.201 0.975
Table 4: GPT-4o Performance - Part 1
Game M.Ratio Win% Score Eff.
Investdie 69.6 0.0 0.219 0.982
Islands 36.1 0.0 0.127 0.361
Jaws 50.1 0.0 0.174 0.244
Killbillvol1 12.6 0.0 0.277 0.000
Labyrinth 39.3 0.0 0.206 0.632
Labyrinthdual 0.0 0.0 0.135 0.000
Lasers 56.3 0.0 0.183 0.221
Lasers2 43.8 0.0 0.222 0.183
Lemmings 28.6 0.0 0.334 0.000
Mirrors 22.2 0.0 0.184 0.000
Missilecommand 62.6 40.0 0.306 0.000
Modality 15.9 20.0 0.330 0.000
Overload 28.4 0.0 0.323 0.000
Pacman 0.0 0.0 0.048 0.000
Painter 42.1 60.0 0.429 0.000
Plants 93.1 0.0 0.318 0.000
Plaqueattack 84.1 20.0 0.345 0.000
Pokemon 48.5 80.0 0.462 0.000
Portals 83.7 0.0 0.253 0.000
Racebet 39.6 0.0 0.153 0.000
Racebet2 51.5 20.0 0.235 0.000
Realportals 29.6 0.0 0.115 0.000
Realsokoban 34.7 0.0 0.372 0.000
Rivers 23.5 0.0 0.250 0.000
Roadfighter 83.9 0.0 0.251 0.000
Roguelike 38.8 0.0 0.199 0.000
Run 72.6 0.0 0.231 0.000
Seaquest 84.8 0.0 0.283 0.000
Sheriff 76.4 0.0 0.247 0.000
Shipwreck 37.4 40.0 0.296 0.000
Sistersavior 51.0 50.0 0.297 0.000
Sokoban 44.7 0.0 0.398 0.000
Solarfox 50.9 0.0 0.177 0.000
Superman 71.2 0.0 0.241 0.000
Surround 22.1 100.0 0.349 0.000
Survivezombies 77.1 20.0 0.324 0.000
Tercio 9.3 0.0 0.183 0.000
Thecitadel 39.0 0.0 0.383 0.000
Themole 56.4 0.0 0.257 0.000
Theshepherd 25.0 0.0 0.348 0.000
Thesnowman 36.2 0.0 0.327 0.000
Trappedhero 0.0 0.0 0.185 0.000
Treasurekeeper 39.4 0.0 0.148 0.000
Vortex 51.2 0.0 0.289 0.000
Waferthinmints 64.2 0.0 0.205 0.000
Waferthinmin… 35.3 100.0 0.376 0.000
Waitforbreak… 13.3 0.0 0.075 0.000
Watergame 39.1 0.0 0.138 0.000
Waterpuzzle 0.0 0.0 0.222 0.000
Waves 70.2 0.0 0.240 0.000
Whackamole 54.3 20.0 0.252 0.000
Wildgunman 71.3 0.0 0.233 0.000
Witnessprote… 98.8 0.0 0.352 0.000
Witnessprote… 73.2 0.0 0.241 0.000
Wrapsokoban 29.2 0.0 0.359 0.000
X-Racer 94.5 0.0 0.297 0.078
Zelda 9.7 0.0 0.083 0.000
Zenpuzzle 31.6 0.0 0.364 0.000
Table 5: GPT-4o Performance - Part 2
Game Without tagger With tagger
aliens 0.00 0.08
boulderdash 0.00 0.00
escape 0.00 0.04
realsokoban 0.00 0.00
sokoban 0.00 0.00
zelda 0.00 0.00
Table 6: Winrate without coordinate tagger vs. with coordinate tagger of ChatGPT-4o-mini. The significance of any improvement is judged using Fisher’s exact test over 25 episodes per game. None of the improvements were statistically significant.
To further examine the generalization ability of GPT-4o-mini, we report detailed results across a broader set of games in the GVGAI-LLM benchmark. Tables 4 and 5 present the model’s performance on 100+ environments, covering meaningful step ratio, win rate, overall behavioral score, and step efficiency. While the model demonstrates high engagement in many environments (e.g., Flower, Infection, Butterflies), it achieves consistently low win rates, particularly in puzzle and planning-heavy games.
We also assess the effect of coordinate-based spatial grounding. Table 6 compares win rates with and without coordinate taggers across six spatially demanding games. Although slight improvements are observed in aliens and escape, none reach statistical significance under Fisher’s exact test, suggesting that coordinate tagging alone is insufficient to resolve core spatial reasoning limitations.
B. Zero-Shot Prompt Example
We design a structured prompt to support zero-shot reasoning in symbolic games. The prompt contains multiple components, including natural language descriptions of the game rules, mechanics analysis, available actions, strategy advice, and a fully verbalised game state. An excerpt of the prompt is shown below (abridged for space):
User
=== Game Rules ===
Game rules in natural language:
# Game Analysis
Genre: This is a puzzle-adventure game with elements of strategy and item collection.
Mechanics:
1. Sprites: The game includes avatars, walls, doors, gems, keys, and exits.
2. Transformation and Interaction: The avatar collects keys, avoids enemies, and triggers effects via sprite interactions.
3. Win/Loss: Reaching the exit wins the game. Losing occurs if the avatar is destroyed.
Strategy Suggestions:
- Collect keys to unlock doors and reach the exit.
- Avoid enemies (gems) unless necessary.
- Use missile keys strategically.
=== Available Actions ===
0: ACTION_NIL
1: ACTION_LEFT
2: ACTION_RIGHT
3: ACTION_DOWN
4: ACTION_UP
=== Important Mechanics Notice ===
Some directional actions may rotate the avatar without movement. Repeating the direction may be needed. Avoid null actions. Interpret the state carefully and act meaningfully.
=== Sprite Mapping ===
avatar $\rightarrow$ ’a’
background floor $\rightarrow$ ’.’
exit $\rightarrow$ ’e’
diamond $\rightarrow$ ’%’
wall $\rightarrow$ ’b’
dirt $\rightarrow$ ’&’
enemy $\rightarrow$ ’\$’
=== Current State ===
bbbbbbbbbbbbbbbbbbbbbbbbbb
b&@@&&&&&@@@........a....b
b&%&&b@b&&&%&&.&&&@@&&&bbb
Each line shows entity at (row, col).
row=1, col=3 $\rightarrow$ a
row=2, col=7 $\rightarrow$ % (diamond)
row=2, col=8 $\rightarrow$ & (dirt)
row=3, col=2 $\rightarrow$ % (diamond)
row=1, col=14 $\rightarrow$ \$ (enemy)
The agent is instructed to output a response in the format: \\ Action:<action number>
It also provides a one-line justification of the action in relation to the current strategy.
This structured prompt enables the LLM to (1) understand symbolic state representations, (2) reason over spatial layout and interaction rules, and (3) produce meaningful, goal-aligned actions in a zero-shot setting.
C. Decision Time Comparison Between LLM Agents and MCTS
Table 7 reports the per-move latency for all LLM agents evaluated in GVGAI-LLM. For models served via the Portkey API, we use the median (p50) latency statistics collected over the past month. For DeepSeek models, latency is measured as the average response time over 10 representative game prompts. In contrast, MCTS is not an API-based model but a symbolic planning algorithm that directly simulates game rules; its decision time is set to the standard per-move time budget (0.04s) defined in the original GVGAI competition.
The results reveal that LLM agents are two to three orders of magnitude slower than symbolic search methods. While real-time playability is not the focus of our benchmark, this latency gap highlights the computational inefficiency of current LLM-based agents compared to classical planning algorithms.
Agent Model Time/Move
Gemini-2.5-pro* LLM 27.82
Gemini-2.0-flash-exp* LLM 1.05
GPT-o3-mini* LLM 35.94
GPT-4o-mini* LLM 2.37
DeepSeek Reasoner LLM 166.00
DeepSeek Chat LLM 1.80
MCTS Search Algorithm 0.04
Table 7: Estimated decision time per move (in seconds) for LLM agents and MCTS.
* Served via Portkey API; Median (p50) latency from Portkey API over the past month.
Latency averaged over prompts from 10 GVGAI games.
Monte Carlo Tree Search (MCTS) agent using symbolic rule simulation; 0.04s is the per-move time budget defined in the original GVGAI competition settings.
D. RL Agent Configuration Details
To compare LLM agents with reinforcement learning (RL) baselines, we trained PPO and DQN agents using standard configurations from the Stable Baselines3 library, with minimal adjustments for symbolic grid-based input. Specifically, both policies employ a custom feature extractor (PaddedEmbeddingCNN) designed to handle discrete int8 observations. All hyperparameters are fixed and shared across games and levels. No tuning was performed, in order to ensure a fair comparison with LLM agents, which are also evaluated in a zero-shot manner.
PPO Configuration (used without tuning)
  • Policy: MlpPolicy with custom extractor (PaddedEmbeddingCNN)
  • Embedding dimension: 32
  • Feature dimension: 256
  • MLP architecture:
    • $\pi$-network: [128, 64]
    • Value-network: [128, 64]
  • Learning rate: 3e-4
  • n_steps: 2048
  • Batch size: 256
  • n_epochs: 10
  • Discount factor ($\gamma$): 0.99
  • GAE lambda: 0.95
  • Entropy coefficient: 0.01
  • Rollout buffer: Int8RolloutBuffer
  • Device: RTX4090
DQN Configuration (used without tuning)
  • Policy: MlpPolicy with custom extractor (PaddedEmbeddingCNN)
  • Embedding dimension: 32
  • Feature dimension: 256
  • Q-network architecture: [128, 64]
  • Learning rate: 3e-4
  • Replay buffer: Int8ReplayBuffer
  • Buffer size: $\min(1{,}000{,}000,\ \text{total timesteps})$
  • Batch size: 256
  • Discount factor ($\gamma$): 0.99
  • Exploration:
    • Initial epsilon: 1.0
    • Final epsilon: 0.02
    • Exploration fraction: 0.1
  • Target update interval: 1000
  • Device: RTX4090
Evaluation Protocol:
All agents, including reinforcement learning (PPO, DQN), search-based (MCTS), and language model agents, evaluated on six GVGAI games: zelda, aliens, boulderdash, escape, sokoban, and realsokoban are under the same experimental protocol. Each game contains 5 levels, resulting in 30 unique game-level combinations. For every level, each agent is evaluated across 5 independent episodes. All reported metrics (e.g., win rate, total reward, number of steps) are averaged over these 5 runs per level, ensuring consistent comparison across models and methods. Heatmaps for LLM win rate are presented separately on each level.